知的財産論 2 単位 (選択)

Introduction to Intellectual Property  藤井章夫 · 客員教授，矢野幹雄 · 非常勤講師，飯田昭夫 · 非常勤講師，山内康伸 · 非常勤講師，京和尚 · 非常勤講師

【授業目的】知的所有権制度を理解し，知的所有権の保護と制度の活用の重要性 ; を各種の事例を基に理解するとともにその活用法の基礎を修得する． ; 【授業概要】科学技術創造立国を目指す我が国において，知的所有権の保護と制 ; 度の活用が，行政 · 産業界 · 大学 · 研究所にとって不可欠であるとの共通認識 ; を持つ必要性を説くとともに，知的所有権制度の概要と，その活用法を，企業 ; の特許戦略，特許侵害事件などの事例を紹介しつつ講義し，学生が将来，企 ; 業 · 大学 · 研究所などへ進んだ場合に知っておくべき知的所有権に関する基礎 ; 知識の修得をはかる． ; 【授業形式】講義 ; 【キーワード】知的財産，特許法，知的財産の保護と活用 ; 【先行科目】 [先行科目] ; 【関連科目】『知的財産論』 (0.5) ; 【履修要件】特になし ; 【履修上の注意】延べ 4 日間の集中講義となり，各集中講義には全時間の出席を要 ; する． ; 【到達目標】 ; 1. 知的財産権の概念についての理解を深める． ; 2. 特許法，商標法，意匠法，著作権法について理解する. ; 3. 知的財産の活用法の基礎について理解する. ; 【授業計画】 ; 1. 知的所有権とは (藤井) ; 2. 知的所有権制度の概要 (特許�商標等) (藤井) ; 3. 知的所有権制度の概要 (意匠�著作権制度等) (藤井) ; 4. 特許発明と特許権侵害 (藤井) ; 5. 特許発明と特許権侵害 (事例) (藤井) ; 6. 知的所有権の管理とその活用 (企業�大学�研究所) (藤井) ; 7. 今後の研究開発と知的所有権のあり方 (藤井) ; 8. レポート課題・書式・提出方法の説明等 (藤井) ; 9. 特許明細書の読み方 -特許は明細書を中心に動く- (山内) ; 10. 技術者が知っておくべき実務知識 -特許法上の留意事項- (山内) ; 11. 技術開発の現場における知的財産 (京和) ; 12. 研究と特許権侵害 (矢野) ; 13. 知的財産の利用と活用 (種苗法による育成者権含む) (飯田)

14. 知的財産の利用と活用 (種苗法による育成者権含む) (飯田)

15. 知的財産の利用と活用 (種苗法による育成者権含む) (飯田)

16. レポート課題・書式・提出方法の説明等 (教務委員長) 【成績評価基準】到達目標が各々達成されているかを試験及びレポートで評価し，

平均で 60%以上であれば合格とする．【教科書】講師が作成したテキスト等を用いる．【参考書】中山信弘著「知的所有権」日刊工業新聞社, 通産省特許庁編「これか

らは日本も知的創造時代」通商産業調査会【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216755 【対象学生】開講コース学生のみ履修可能【連絡先】

⇒ 学務係 (オフィスアワー: 月曜から金曜:8時30分∼ 17時15分)

Introduction to Intellectual Property 2 units (selection)

Akio Fujii · Visiting Professor , Mikio Yano · Part-time Lecturer , Akio Iida · Part-time Lecturer , Yasunobu Yamauchi · Part-time Lecturer

, 名姓 · Part-time Lecturer

Target⟩ Understanding the basic systems and importance of protection and practical

use of intellectural property based on some examples. Outline⟩ It is necessary for a working staff in a company, a university, etc., in a

future.It is introduced the importance to protect and practical use of intellectural property at governments , industrial circles, univesities and institutes in Japan using potential on scientific techniques. This lecture is treated the basic systems of protection and practical use of intellectural property particularly based on strategy and invasion of patent in an organization such as company, university, etc.

Style⟩ Lecture �������⟩ intellectual property, patent law, protect and use of intellectural

property Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ “Introduction to Intellectual Property”(0.5) Requirement⟩ No requierement Notice⟩ It is necessary to be present at Four days concentrated lecture.

Goal⟩

1. Understanding concept of intellectural property right.

2. Undestand the Patent Act, the Trademark Act,the Design Act and the Copyright Act.

3. Understanding on basical practical use of intellectural property.

Schedule⟩ 1. 知的所有権とは (藤井)

2. 知的所有権制度の概要 (特許�商標等) (藤井)

3. 知的所有権制度の概要 (意匠�著作権制度等) (藤井)

4. 特許発明と特許権侵害 (藤井)

5. 特許発明と特許権侵害 (事例) (藤井)

6. 知的所有権の管理とその活用 (企業�大学�研究所) (藤井)

7. 今後の研究開発と知的所有権のあり方 (藤井)

8. レポート課題・書式・提出方法の説明等 (藤井)

9. 特許明細書の読み方 -特許は明細書を中心に動く- (山内)

10. 技術者が知っておくべき実務知識 -特許法上の留意事項- (山内)

11. 技術開発の現場における知的財産 (京和)

12. 研究と特許権侵害 (矢野)

13. 知的財産の利用と活用 (種苗法による育成者権含む) (飯田)

14. 知的財産の利用と活用 (種苗法による育成者権含む) (飯田)

15. 知的財産の利用と活用 (種苗法による育成者権含む) (飯田)

16. レポート課題・書式・提出方法の説明等 (教務委員長)

Evaluation Criteria⟩ Passing mark is over 60% on the examinations and the

reports for each reaching level, respectively. Textbook⟩ To be used textbooks made by lecturers. Reference⟩ Nobuhiro Nakayama:Titeki Shoyuken, Nikkan Kohgyo Shinbunnsha

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216755

Student⟩ Able to be taken by only specified class(es)

Contact⟩ ⇒ School Affairs (Office Hour: 月曜から金曜:8時30分∼ 17時15分)